imqwerty (imqwerty):

u got 2 pills in ur pocket one red one blue and u r blind nd u wanna eat half of each how will u do so??? :))

OpenStudy (hayleymeyer):

this question.. haha xD

OpenStudy (anonymous):

blue

OpenStudy (anonymous):

lol ^_^

Parth (parthkohli):

I'll just eat half of both pills.

imqwerty (imqwerty):

how will u do so @ParthKohli

OpenStudy (texaschic101):

just break each pill in half

OpenStudy (anonymous):

Get tired of try to break them in half and eat them both whole

OpenStudy (anonymous):

Simply put both pills in your hands, and break each pill in half.

OpenStudy (kyanthedoodle):

I'd take the purple pill. Purple always wins.

pooja195 (pooja195):

i'll take both

Parth (parthkohli):

Also, why do I just want to each half of each pill? Are they medicinal pills? If I want to eat them, why don't I just completely eat them?

OpenStudy (texaschic101):

if the pills are scored -- there is a line in the middle-- makes it easy to break.

OpenStudy (texaschic101):

now you would be in a bit of a jam if they were capsules...lol

OpenStudy (kyanthedoodle):

Waaait a minute. If the pills are in my pocket, and I'm blind, how did they get there in the first place? Did some creepy guy put them there?

OpenStudy (hayleymeyer):

omgg, this is funny.

Parth (parthkohli):

I'll do so using my mouth.

OpenStudy (texaschic101):

you can be blind and still find your pockets to put pills in..lol

imqwerty (imqwerty):

lol XD

OpenStudy (kyanthedoodle):

Then that means that I can be blind, and take them back out! Aha! I have figured it out!

OpenStudy (texaschic101):

ROFL...

OpenStudy (anonymous):

:O But what type of pockets are they????

OpenStudy (anonymous):

Like do they have zippers?

OpenStudy (texaschic101):

blind people can undo buttons as well

OpenStudy (texaschic101):

or pull zippers....they are not invalids

OpenStudy (anonymous):

Or like a lock

OpenStudy (anonymous):

if I'm blind, how I am i reading this question in the first place? :O

Parth (parthkohli):

Blind people can also break pills in half.

OpenStudy (kyanthedoodle):

Ooooooooooooooooooooooooooooooooooooooh snap

OpenStudy (anonymous):

and he can't see what the combination is.

imqwerty (imqwerty):

the pockets r sealed with a password nd u need to crack it with the help of this code that u have to solve the get the password @brandonford7 -- value of log(iota)

Parth (parthkohli):

Who cares - I want to eat half of both pills.

OpenStudy (texaschic101):

lol...a combination lock pocket...

OpenStudy (hayleymeyer):

interestinggg <333

OpenStudy (anonymous):

hi <3

Parth (parthkohli):

\[i = \cos \pi/2 + i \sin \pi/2 = e^{i\pi/2}\]Just take the log.

OpenStudy (anonymous):

How would I solve it though if I'm blind

OpenStudy (texaschic101):

forget the code ? Just cut a whole in the pocket and get your pills

OpenStudy (anonymous):

You would have to be carful to not cut your self though

imqwerty (imqwerty):

how is i=cos(pi/2) @ParthKohli ( ͡° ͜ʖ ͡°)

imqwerty (imqwerty):

@texaschic101 lol

Parth (parthkohli):

I said \(i = \cos(\pi/2) + i \sin(\pi/2) = e^{i\pi/2}\)

OpenStudy (texaschic101):

now if you had a suit of armor on...well, then you can't cut your pocket

imqwerty (imqwerty):

sorry i didn't read the whole lol @ParthKohli

Parth (parthkohli):

\[\log(i) = i\pi/2 \]

OpenStudy (kyanthedoodle):

Dude. This is too much work. I'd rather just go to the store and get new pills.

imqwerty (imqwerty):

well how will u cut a hole in ur pocket ?

OpenStudy (anonymous):

So there's a blind man wearing armor trying to cut his pills out of his locked pocket? lol

Parth (parthkohli):

Hey, oops. I was wrong.

OpenStudy (texaschic101):

ROFL

TheSmartOne (thesmartone):

I wouldn't eat it in the first place ;p

Parth (parthkohli):

\[\log(i) = i(\pi/2 + 2n \pi)\]

imqwerty (imqwerty):

well u got the password but how will u type the answer or like how will u open the lock after u know the answer

Parth (parthkohli):

With the key.

imqwerty (imqwerty):

lol

OpenStudy (anonymous):

number pads?

TheSmartOne (thesmartone):



OpenStudy (texaschic101):

its a braille lock

Parth (parthkohli):

I'll type my answer as well as open the locks with keys. Hehe.

imqwerty (imqwerty):

XD

imqwerty (imqwerty):

from where will u buy such a fancy pant??

Parth (parthkohli):

Rule 34.

TheSmartOne (thesmartone):

From the blind man

imqwerty (imqwerty):

smart answer by @TheSmartOne

Parth (parthkohli):

log(i) is not well-defined. Just thought I'd get that through.

TheSmartOne (thesmartone):

Certainly, we all can't be that one blind man :p

TheSmartOne (thesmartone):

Or you just go and help the blind man eat this mysterious pills x'D

OpenStudy (empty):

I will just eat both pills since eating two whole pills implies I ate both half pills.

OpenStudy (texaschic101):

overdose

OpenStudy (empty):

Good.

OpenStudy (texaschic101):

now your a dead blind man

imqwerty (imqwerty):

cool

OpenStudy (kyanthedoodle):

Whatever. My life as a blind person wasn't that great anyway. XD

OpenStudy (empty):

I deserve it for not being able to figure this out, this is probably how I got blinded in the first place; by being a complete idiot.

imqwerty (imqwerty):

who all have read 'The story of my life' by Helen Keller

OpenStudy (texaschic101):

well...I saw the movie...there is no way I could have been her teacher

imqwerty (imqwerty):

ok nw u need to be real smart to solve this question - u got a gum stuck in your hair how will u get rid of it??

imqwerty (imqwerty):

without cutting your hair

OpenStudy (anonymous):

Cut off my head :)

imqwerty (imqwerty):

nd without dying too

OpenStudy (anonymous):

pull my hair out

TheSmartOne (thesmartone):

Or just eat one pill

OpenStudy (anonymous):

Darn it =(

TheSmartOne (thesmartone):

1 pill = 1/2 pill + 1/2 pill

TheSmartOne (thesmartone):

\(\color{blue}{\text{Originally Posted by}}\) @imqwerty ok nw u need to be real smart to solve this question - u got a gum stuck in your hair how will u get rid of it?? \(\color{blue}{\text{End of Quote}}\) peanut butter, always works :p

OpenStudy (triciaal):

if blind how do you know the pills are red and blue? one in each hand and only take half of each

OpenStudy (anonymous):

I'd go with putting my hair in hot water and removing the gum

pooja195 (pooja195):

LOL

TheSmartOne (thesmartone):

triciaal has a point!

imqwerty (imqwerty):

haha i forgot that @triciaal

imqwerty (imqwerty):

nw give ur answer smart people

TheSmartOne (thesmartone):

answer to life?

imqwerty (imqwerty):

answer to my question

OpenStudy (anonymous):

Is putting your hair in hot water not a valid option?

imqwerty (imqwerty):

yes u got it right @LegendarySadist